Soybean meal is 14% protein, cornmeal is 7% protein. How many pounds of each should be mixed together to get a 280 pound mixture that is 8% protein

How many pounds of the cornmeal should be in the mixture?
___pounds?

Answer by mananth(16949)

Soybean meal is 14% protein, cornmeal is 7% protein. How many pounds of each should be mixed together to get a 280 pound mixture that is 8% protein
percent ---------------- quantity
Soyabean 14----------------x
Corn 7----------------280-x
Total 8.00%---------------- 280
14x+7(280-x)=8* 280
14x+1960 -7x=2240
14x-7x= 2240-1960
7x=280
/7
x=40 soyabean pounds
corn = 280 -40 = 240 pounds


240